OpenStudy (anonymous):

sin20.sin40.sin60.sin80 is??

OpenStudy (anonymous):

sin(20) sin(40) sin (60) sin (80) substitute sin(60) = √3 /2 √3/2 [ sin(20) sin(40) sin(80) ] = (√3/2) sin(20) [ sin(40) sin(80) ] use the formula sin A sin B = (1/2) [ cos(A - B) - cos(A + B) ] = √3/2 sin(20) (1/2)[ cos(40) - cos(120) ] = √3/4 sin(20) [ cos(40) + cos(60) ] = √3/4 sin(20) [ cos(40) + 1/2 ] = √3/4 sin(20)cos(40) + (√3/8) sin(20) use the formula sin A cos B = 1/2 [ sin(A + B) + sin(A - B) ] = (√3/4)(1/2) [ sin(60) + sin(-20) ]+ (√3/8)sin(20) = (√3/8) [ (√3 / 2) - sin(20) ]+ (√3/8)sin(20) = 3/16 - (√3/8)sin(20) + (√3/8)sin(20) = 3/16